OpenStudy (anonymous):

Darian is buying an entertainment center for his living room wall. The tag on his first choice shows that the furniture is forty inches wide and ninety inches high. Darian has a space for it that is 5 feet wide and 7 feet high. Will his first choice fit? Explain.

OpenStudy (anonymous):

use unit analysis to convert inches to feet, then calc. the area and see if fits.

OpenStudy (anonymous):

(40")(1'/12")=3&(1/3)' (90")(1'/12")=7.5'

OpenStudy (anonymous):

(10/3)'*(7.5)' = 25ft^2 5'*7'=35ft^2 entertainment center < Darian's space

OpenStudy (anonymous):

3&(1/3) ft = (10/3) ft

OpenStudy (anonymous):

Do you understand?